# GCoT: Chain-of-Thought Prompt Learning for Graphs

### v1

@@ 4.2. Chain-of-Thought Prompting
where $w^{1},w^{2},\cdots,w^{L}$ are learnable parameters. The thought $\mathbf{T}_{k}$ is then used to guide the the next inference step.
Thought conditioned prompt learning The thought generated in the previous inference step stores hierarchical structural knowledge, which guides the learning in the subsequent step. Since different nodes may exhibit distinct characteristics with respect to the downstream task, rather than learn all node’s representation in the same manner, the pre-trained model should be steered to learn node-specific patterns. To this end, we propose leveraging a condition-net (Zhou et al., 2022; Yu et al., 2024b, 2025) to generate node-specific prompts, guiding node-specific adaptation. Specifically, conditioned on the thought from the previous step, $\mathbf{T}_{k}$, the condition-net generates a series of thought prompts as follows:
(7) $\mathbf{P}_{k}=\mathtt{CondNet}(\mathbf{T}_{k};\phi),$
where $\mathtt{CondNet}$ is the condition-net parameterized by $\phi$. Note that condition-net is a lightweight hypernetwork (Ha et al., 2022), where our condition-net $\mathtt{CondNet}$ functions as an auxiliary network to generate the prompts conditioned on the thoughts parameter-efficiently. In our implementation, we utilize a simple multi-layer perceptron (MLP) with a bottleneck structure for improved efficiency (Wu and Lee, 2018). The $i$-th row of $\mathbf{P}_{k}$ represents a unique node-specific prompt vector $\mathbf{p}_{i,k}$ for node $v_{i}$, which is derived from the thought $\mathbf{T}_{k}$. This prompt vector then guides the next inference step by modifying the node features as follows:
@@
where $\odot$ denotes element-wise multiplication, and $\mathbf{X}_{k+1}$ represents the input to the pre-trained graph encoder in the $k+1$ step. After repeating $K$ inference steps, the final step outputs the nodes embeddings $\mathbf{H}_{K}$.
@@
Prompt tuning For initial prompts, all standard graph prompting methods can be applied within our framework. In our experiments, we adhere to GPF+ (Fang et al., 2024) to generate initial prompts. Specifically, we train $N$ bias prompts $\{\mathbf{p}_{\text{bias}}^{1},\mathbf{p}_{\text{bias}}^{2},\ldots,\mathbf{p}_{\text{bias}}^{N}\}$, and leverage attention-based aggregation to generate node-specific prompts. However, unlike GPF+, which uses these prompts to modify graph features, we use them to modify the final embeddings. Specifically, the task prompt for node $i$ is computed as:
(9) $\mathbf{p}_{\text{init}}^{i}=\sum_{j=1}^{N}\alpha_{i,j}\mathbf{p}_{\text{bias}}^{j},\quad\text{where}\quad\alpha_{i,j}=\frac{\exp\big{(}\mathbf{a}^{j}\mathbf{h}_{K}^{i}\big{)}}{\sum_{l=1}^{N}\exp\big{(}\mathbf{a}^{l}\mathbf{h}_{K}^{i}\big{)}}.$
@@
Here, $\mathbf{h}_{K}^{i}$ denotes the $i$-th row of the final embedding matrix $\mathbf{H}_{K}$, which serves as the embedding for node $i$, and $\mathbf{A}=\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{N}\}$ are $N$ learnable linear projection vectors. The generated initial prompts form the matrix $\mathbf{P}_{\text{init}}$, which is then used to modify the final output embeddings as follows:
(10) $\mathbf{\tilde{H}}=\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}.$
@@
Note that for clarity and to maintain a unified representation of the various task prompt mechanisms, we use $\mathbf{P}_{\text{init}}$ to denote the trainable parameters associated with these mechanisms throughout the remainder of this paper.
Finally, for a given task with a labeled training set
$\mathcal{D}_{t}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\},$
@@
where each $x_{i}$ represents either a node or a graph and $y_{i}\in Y$ denotes its corresponding class label, the downstream loss function is defined as:
(11) $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})=-\sum_{(x_{i},y_{i})\in\mathcal{D}_{t}}\ln\frac{\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{y_{i}}\right)/\tau\right)}{\sum_{c\in Y}\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{c}/\right)\tau\right)},$
@@
where $\mathbf{\tilde{h}}_{x_{i}}$ represents the final embedding of a node $v$ or a graph $G$. Specifically, for node classification, $\mathbf{\tilde{h}}_{v}$ corresponds to a row in the answer matrix $\mathbf{\tilde{H}}$, while for graph classification, we compute the graph embedding as
$\mathbf{\tilde{h}}_{G}=\sum_{v\in V}\mathbf{\tilde{h}}_{v},$
@@
incorporating an additional graph readout step. The prototype embedding for each class $c$, denoted as $\mathbf{\bar{h}}_{c}$, is obtained by averaging the embeddings of all labeled nodes or graphs belonging to that class. During prompt tuning, only the task prompts and lightweight parameters of the condition network ($\phi$) are updated, while the pre-trained GNN weights remain frozen. This parameter-efficient design makes our approach well-suited for few-shot learning, where the training set $\mathcal{D}_{t}$ contains only a limited number of labeled examples.
@@ 4.3. Algorithm and Complexity Analysis
Algorithm. We detail the main steps for Chain-of-Thought graph prompting in Algorithm 1. In lines 3–12, we iterate through $K$ inference steps during the downstream adaptation phase while keeping the pre-trained weights $\Theta_{0}$ frozen. Specifically, in lines 4–7, we generate the thought for the $k$-th inference step by fusing the output embeddings from each layer of the pre-trained graph encoder. In lines 8–12, we leverage this thought to generate node-specific prompts that guide the subsequent inference step to capture node-specific patterns. In lines 13–14, we employ a standard task prompt to modify the output embedding of the last inference step. Note that, based on the task prompt mechanism described in the original paper, these prompts can also be applied to modify the input features at the first inference step. Finally, in lines 15–17, we update the embeddings for the prototypical nodes/graphs based on the few-shot labeled data.
@@
1:Pre-trained graph encoder with parameters $\Theta_{0}$.
2:Optimized parameters $\phi$ of condition-net, and task prompt $\mathbf{P}_{\text{init}}$.
3:while not converged do
4:    $\phi_{i}\leftarrow$ initialization
@@
8:         /* Encoding by the pre-trained graph encoder */
9:         $\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{k};\Theta_{0})$
@@
12:         /* Thought prompting */
13:         /* Generate thought prompts by Eq. 7 */
14:         $\mathbf{P}_{k}\leftarrow\mathtt{CondNet}(\mathbf{T}_{k};\phi)$
15:         /* Thought prompts modification by Eq. 8 */
16:         $\mathbf{X}_{k+1}\leftarrow\mathbf{P}_{k}\odot\mathbf{X}$
17:    /* $K$-th inference step */
18:    $\{\mathbf{H}_{K}^{1},\mathbf{H}_{K}^{2},\cdots,\mathbf{H}_{K}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{K};\Theta_{0})$
19:    $\mathbf{H}_{K}\leftarrow\mathbf{H}_{K}^{L}$
20:    /* Initial prompt modification by Eq. 10*/
21:    $\mathbf{\tilde{H}}\leftarrow\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}$
22:    /* Update prototypical instance */
23:    for each class $c$ do
@@
24:         ${\mathbf{\bar{h}}}_{c}\leftarrow\textsc{Average}(\mathbf{\tilde{h}}_{x}$: instance $x$ belongs to class $c$)
25:    /* Optimizing the parameters in condition-net */
26:    Calculate $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$ by Eq. (11)
27:    Update $\phi$ by backpropagating $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$
28:return $\{\phi,\mathbf{P}_{\text{task}}\}$
@@
Complexity analysis. For a downstream graph $G$, we perform $K$ iterative inference steps. Each step comprises two key components: (1) thought generation using a pre-trained GNN (executed for $K$ iterations) and (2) thought-based conditional prompt learning (executed for $K-1$ iterations). The computational complexity of the first component is largely determined by the GNN architecture. In a standard GNN, each node aggregates information from up to $D$ neighboring nodes per layer. Consequently, computing node embeddings over $L$ layers incurs a complexity of $O(D^{L}\cdot|V|)$, where $|V|$ denotes the number of nodes. Aggregating the outputs from all $L$ layers adds an extra complexity of $O(L\cdot|V|)$. Since the thought generation process is repeated $K$ times, its overall complexity is:$O(K\cdot(D^{L}+L\cdot|V|))$. The second component, thought-based conditional prompt learning, refines the learning process by leveraging the thought representation from the previous step. This component consists of two stages: prompt generation and prompt tuning. In the prompt generation stage, the thought is fed into the condition-net, with a complexity of $O(|V|)$. In the prompt tuning stage, each node is adjusted via a prompt vector, also incurring a complexity of $O(|V|)$. Since this process is executed for $K-1$ iterations, the total complexity for this phase is: $O(2(K-1)\cdot|V|)$. Additionally, we employ a standard task prompt to modify node features or embeddings. The complexity of this process depends on the specific prompting mechanism. Here, we assign it a complexity of $O(|V|)$, which is common among graph prompting methods (Liu et al., 2023; Fang et al., 2024). In summary, the overall computational complexity of GCoT is: $O((D^{L}+L+2(K-1)+1)\cdot|V|)$.
4.4. Why GCoT Works
In GCoT, we adapt the following mechanism. (1) All three approaches leverage a universal task template to unify pre-training and downstream tasks, ensuring that GCoT can efficiently adapt to different downstream tasks, particularly in few-shot settings. (2) The chain-of-thought mechanism enables GCoT to perform multiple inference steps, thereby refining the final answer more effectively. (3) The thoughts in GCoT fuse hierarchical topological knowledge from graphs, allowing it to better capture structural information. (4) Conditioned on the thought from the previous step, GCoT generates a series of node-specific prompts that efficiently capture fine-grained node characteristics, thus effectively refining answer step by step. The combination of above mechanisms ensures the effectiveness of GCoT.
@@ 5. Experiments
indicates the downstream task category for each dataset: ”N” corresponds to node classification, while ”G” signifies graph classification.
Table 2. Accuracy (%) evaluation on node and graph classification.
@@
DGI/InfoGraph 54.11±9.60 45.00±9.19 57.98±11.36 64.48±10.04 53.17±17.29 53.82±14.19 49.33±15.11 52.51±10.29
GraphCL 51.96±9.43 43.12±9.61 61.18±11.60 63.93±10.21 54.92±17.09 53.81±14.21 49.73±14.66 53.81±8.97
ProG 51.35±13.92 35.08±8.63 64.07±14.74 68.85±9.99 51.99±4.50 53.45±15.01 53.52±11.97 52.73±6.57
GPF 57.35±13.87 43.04±8.85 57.61±11.26 65.26±9.89 56.55±13.95 54.16±14.07 48.65±13.96 53.05±7.62
GPF+ 64.91±15.49 42.28±8.71 57.85±10.86 69.80±10.84 56.81±12.93 55.24±13.29 50.83±19.74 54.58±8.70
GraphPrompt 54.56±12.56 46.23±8.70 64.85±10.72 68.99±9.28 55.44±12.56 54.34±14.77 54.59±10.52 53.80±7.93
GCoT 66.09±15.17 47.77±10.93 70.57±17.10 77.51±11.76 58.75±15.42 56.26±15.52 58.03±23.44 56.24±8.60

### v2

@@ 4.2.1. Inference step
where $W=\{w^{l}\in\mathbb{R}:1\leq l\leq L\}$ are learnable parameters. The resulting thought, $\mathbf{T}_{k}$, captures the current working state of the graph encoder, storing multi-layer structural knowledge. In particular, the $i$-th row of $\mathbf{T}_{k}$ reflects the thought pertaining specifically to node $v_{i}$.
Thought-conditioned prompt learning. The thought constructed in Eq. (6) is then used for prompt generation, in order to further guide the next inference step. Moreover, since different nodes may exhibit distinct characteristics with respect to the downstream task, it is beneficial to adapt the pre-trained model to node-specific patterns. Thus, rather than employing a single prompt for all nodes, we propose to leverage a conditional network (condition-net) (Zhou et al., 2022; Yu et al., 2025b, c) to generate node-specific prompts and facilitate node-specific adaptation. Specifically, conditioned on the current thought $\mathbf{T}_{k}$, the condition-net generates a series of prompts, collectively represented as $\mathbf{P}_{k}\in\mathbb{R}^{|V|\times d}$, as follows.
(7) $\mathbf{P}_{k}=\textsc{CondNet}(\mathbf{T}_{k};\phi),$
where CondNet is the condition-net parameterized by $\phi$. The condition-net can be viewed as a hypernetwork (Ha et al., 2022)—a lightweight auxiliary network such as a multi-layer perceptron (MLP)—that generates a distinct prompt for each node from the thought, yet avoids parameterizing a separate prompt vector for each node. In particular, the $i$-th row of $\mathbf{P}_{k}$ represents a unique node-specific prompt vector $\mathbf{p}_{k,i}$ for node $v_{i}\in V$. Subsequently, the prompt vectors are used to modify the node features of the query graph for prompt-based inference in the next (i.e., $k+1$-th) step, as follows.
@@
where $\odot$ denotes element-wise multiplication, and $\mathbf{X}_{k+1}$ represents the input to the pre-trained graph encoder in the $k+1$-th step.
4.2.2. Standard prompt learning
To align the objectives of pre-training and downstream tasks, we also leverage a standard prompt following prior work, which typically modifies node features (Sun et al., 2022; Fang et al., 2024; Yu et al., 2024c) or embeddings (Liu et al., 2023; Yu et al., 2024c). In particular, we adopt GPF+ (Fang et al., 2024) to generate the standard prompts. However, we emphasize that our framework is compatible with any standard graph prompting technique. As demonstrated in Sect. 5.6, our CoT-style prompting can be combined with several prevailing graph prompting approaches to further enhance their performance.
@@ 4.2.2. Standard prompt learning
Specifically, we train $N$ bias prompts $\{\mathbf{p}_{\text{bias}}^{1},\ldots,\mathbf{p}_{\text{bias}}^{N}\}$ and leverage attention-based aggregation to generate node-specific prompts. While GPF+ (Fang et al., 2024) applies these prompts to graph features $\mathbf{X}$, we use them to modify the output embeddings $\mathbf{H}_{K}$ after $K$ inference steps. Concretely, the standard prompt for node $v_{i}$ is computed as
(9) $\mathbf{p}_{\text{std},i}=\sum_{j=1}^{N}\alpha_{i,j}\mathbf{p}_{\text{bias}}^{j},\quad\text{where}\ \alpha_{i,j}=\frac{\exp\big{(}\mathbf{a}^{j}\mathbf{h}_{K,i}\big{)}}{\sum_{n=1}^{N}\exp\big{(}\mathbf{a}^{n}\mathbf{h}_{K,i}\big{)}}.$
@@
Here, $\mathbf{h}_{K,i}$ denotes the $i$-th row of the output embedding matrix $\mathbf{H}_{K}$, i.e., the embedding vector of node $v_{i}$ after $K$ inference steps. $\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{N}\}$ are $N$ learnable linear projection vectors. The standard prompts for all nodes are stacked to form the matrix $\mathbf{P}_{\text{std}}$, which is then used to modify the output embeddings $\mathbf{H}_{K}$:
(10) $\mathbf{\tilde{H}}=\mathbf{P}_{\text{std}}\odot\mathbf{H}_{K}.$
We call $\mathbf{\tilde{H}}$ the answer matrix, since it is used by downstream tasks to produce final predictions (or answers) in Sect. 4.3.
@@
Note that for clarity and to maintain a unified representation of various standard prompt mechanisms, we use $\mathbf{P}_{\text{std}}$ to denote the trainable parameters associated with these mechanisms throughout the remainder of this paper.
4.2.3. Summary
In GCoT, we adopt the following mechanisms, which collectively ensure the effectiveness of GCoT. (1) We follow standard graph prompt learning methods (Liu et al., 2023; Fang et al., 2024) to align pre-training and downstream tasks, ensuring that GCoT can efficiently adapt to different downstream tasks even in few-shot settings. (2) The CoT-style prompting performs multiple inference steps, allowing iterative refinement of its prediction for a given downstream task. (3) The thoughts in GCoT fuse hierarchical topological knowledge from graphs, enabling the capture of fine-grained structural information. (4) Conditioned on the thought, GCoT generates a series of node-specific prompts that reflect individualized node characteristics in a parameter-efficient manner.
4.3. Prompt Tuning
Consider a downstream task with a labeled training set
$\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\},$
@@ 4.3. Prompt Tuning
where each $x_{i}$ represents either a node or a graph and $y_{i}\in Y$ denotes its corresponding class label. Subsequently, the loss function of the task is defined as $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})=$
(11) $-\sum_{(x_{i},y_{i})\in\mathcal{D}}\ln\frac{\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\tilde{h}}_{y_{i}}\right)/\tau\right)}{\sum_{c\in Y}\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\tilde{h}}_{c}\right)/\tau\right)},$
@@
where $\mathbf{\tilde{h}}_{x_{i}}$ represents the final embedding of a node $v$ or a graph $G$. Specifically, for node classification, $\mathbf{\tilde{h}}_{v}$ corresponds to a row in the answer matrix $\mathbf{\tilde{H}}$; for graph classification, we apply a readout operation and compute the graph embedding as $\mathbf{\tilde{h}}_{G}=\sum_{v\in V}\mathbf{\tilde{h}}_{v}$. Lastly, $\mathbf{\tilde{h}}_{c}$ denotes the prototype embedding for class $c$, which is obtained by averaging the embeddings of all labeled nodes or graphs belonging to that class.
@@
During prompt tuning, only the weights of thought construction ($W$) and the condition-net ($\phi$), as well as the standard prompts ($\mathbf{P}_{\text{std}}$), are updated, while the pre-trained weights of the graph encoder remain frozen. This parameter-efficient design makes our approach well-suited for few-shot learning, where the training set $\mathcal{D}$ contains only a few labeled examples.
@@ 4.4. Algorithm and Complexity Analysis
1:Pre-trained graph encoder with parameters $\Theta_{0}$, labeled data $\mathcal{D}$ for a given downstream task
2:Optimized parameters $W,\phi,\mathbf{P}_{\text{std}}$.
3:$W,\phi,\mathbf{P}_{\text{std}}\leftarrow$ initialization
4:while not converged do
@@
8:         /* Prompt-based inference */
9:         $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{k};\Theta_{0})$
@@
12:         /* Thought-conditioned prompt generation by Eq. (7) */
13:         $\mathbf{P}_{k}\leftarrow\mathtt{CondNet}(\mathbf{T}_{k};\phi)$
14:         /* Feature modification by Eq. (8) */
15:         $\mathbf{X}_{k+1}\leftarrow\mathbf{P}_{k}\odot\mathbf{X}$
16:    /* Last (i.e., $K$-th) inference step */
17:    $\{\mathbf{H}_{K}^{1},\mathbf{H}_{K}^{2},\cdots,\mathbf{H}_{K}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{K};\Theta_{0})$
18:    $\mathbf{H}_{K}\leftarrow\mathbf{H}_{K}^{L}$
19:    /* Standard prompt modification by Eq. (10) */
20:    $\mathbf{\tilde{H}}\leftarrow\mathbf{P}_{\text{std}}\odot\mathbf{H}_{K}$
21:    /* Update prototypical instance */
22:    for each class $c$ do
@@
23:         ${\mathbf{\tilde{h}}}_{c}\leftarrow\textsc{Mean}(\{\mathbf{\tilde{h}}_{x}$: $x$ is any instance in class $c$ from $\mathcal{D}$})
24:    /* Optimizing the parameters */
25:    Calculate $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})$ by Eq. (11)
26:    Update $W,\phi,\mathbf{P}_{\text{std}}$ by backpropagating $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})$
27:return $\{W,\phi,\mathbf{P}_{\text{std}}\}$
Algorithm. We outline the main steps for GCoT in Algorithm 1. In lines 3–13, we iterate through the first $K-1$ inference steps for a downstream task while keeping the pre-trained weights $\Theta_{0}$ frozen. Specifically, in lines 8–9, we generate the thought for the $k$-th inference step by fusing the embeddings from each layer of the pre-trained graph encoder. In lines 10–11, we leverage this thought to generate node-specific prompts that guide the next inference step. In lines 14–18, we perform the last inference step, employing a standard prompt to modify the output embedding and obtaining the answer matrix. Finally, in lines 19–24, we calculate and backpropagate the downstream loss to update the parameters $W,\phi,\mathbf{P}_{\text{std}}$.
@@
Complexity analysis. For a downstream graph $G$, we perform $K$ inference steps. In each step, we first conduct prompt-based inference, the complexity of which is determined by the architecture of the graph encoder. In a standard GNN, each node aggregates information from up to $D$ neighboring nodes per layer. Consequently, computing node embeddings over $L$ layers results in a complexity of $O(D^{L}|V|)$, where $|V|$ denotes the number of nodes. This is followed by thought generation, which fuses embeddings from all $L$ layers, introducing an additional complexity of $O(L|V|)$. The generated thought is subsequently used for thought-conditioned prompt learning, with a complexity of $O(|V|)$. Therefore, the computational complexity for the $K$ inference steps is $O(K(D^{L}+L)|V|)$. Additionally, we employ a standard prompt to modify node features or embeddings. The complexity of this process depends on the specific prompting mechanism. Here, we assign it a complexity of $O(|V|)$, which is common among graph prompting methods (Liu et al., 2023; Fang et al., 2024). Thus, the overall complexity of GCoT is $O(K(D^{L}+L)|V|)$, or more simply $O(KD^{L}|V|)$, since it is typical that $L\ll D^{L}$.
@@ 5. Experiments
This column indicates the type of downstream task conducted for each dataset: “N” denotes node classification, while “G” denotes graph classification.
Table 2. Accuracy (%) evaluation of node and graph classification.
@@
DGI/InfoGraph 54.11 ±   9.60 45.00 ±   9.19 47.46 ± 12.19 58.89 ± 10.97 53.17 ± 17.29 53.82 ± 14.19 49.33 ± 15.11 52.51 ± 10.29
GraphCL 51.96 ±   9.43 43.12 ±   9.61 46.80 ±   9.04 57.78 ± 11.31 54.92 ± 17.09 53.81 ± 14.21 49.73 ± 14.66 53.81 ±   8.97
ProG 50.59 ± 14.64 43.17 ±   8.49 63.07 ± 11.96 66.50 ±   9.46 51.99 ±   4.50 53.45 ± 15.01 53.52 ± 11.97 52.73 ±   6.57
GPF 57.60 ± 13.88 43.11 ±   8.80 55.63 ± 10.96 65.29 ± 10.07 56.55 ± 13.95 54.16 ± 14.07 48.65 ± 13.96 53.05 ±   7.62
GPF+ 57.42 ± 13.87 43.28 ±   8.82 57.16 ± 10.99 65.07 ± 10.01 56.81 ± 12.93 55.24 ± 13.29 50.83 ± 19.74 54.58 ±   8.70
GraphPrompt 54.25 ±   9.38 45.34 ± 10.53 63.11 ± 10.01 66.62 ±   9.90 55.44 ± 12.56 54.34 ± 14.77 54.59 ± 10.52 53.80 ±   7.93
GCoT 59.67 ± 15.51 46.21 ±  8.78 64.43 ±  9.96 67.16 ± 10.46 58.75 ± 15.42 56.26 ± 15.52 58.03 ± 23.44 56.24 ±  8.60